# Uncertainty in predicting the stochastic gravitational wave background from compact binary coalescences

### v2

@@ IV.4 Combined merger rate uncertainty
To construct the full ensemble of merger rate histories across all classes of compact binaries, we draw random samples from the parameter space defined earlier, which governs the evolution of the merger rate. The sampling ranges and underlying distributions for these parameters are summarized in Table 1. Specifically, we vary the local rate $R_{0}$, the minimum delay time $t_{\mathrm{d,min}}$, the spectral index $\nu$ of the time delay distribution, and exclusively for BBH systems, the metallicity threshold $Z_{\mathrm{th}}$, above which BBH formation is suppressed, and the metallicity spread $\sigma_{Z}$, which characterizes the dispersion around the mean metallicity given by Section IV.3. We adopt a Gaussian prior for $\nu$ (with $\mu_{\nu}=-1$ and $\sigma_{\nu}=0.25$) to reflect results from observations that a spectral index $\nu=-1$ provides a good fit to many compact binary formation models [109, 110, 59]. This choice helps to avoid a bias towards high merger rates that could arise by large negative values of $\nu$, without affecting the overall uncertainty in the combined merger rate uncertainty. It simply ensures a greater representation of merger rates consistent with a time delay distribution with spectral index close to -1. While BNS formation is only weakly sensitive to metallicity, the situation is less clear for NSBH systems. Nevertheless, given the already broad uncertainties in the NSBH merger rate estimates, we choose not to introduce additional uncertainty from metallicity effects. The resulting CBC merger rates are shown in Fig. 11. While the local BBH merger rate is relatively well constrained by direct observations, the rate at redshifts $z>2$ can vary by more than two orders of magnitude depending on the model assumptions. The local rates of BNS and NSBH coalescences are much less constrained.
@@
Figure 11: Ensembles of merger rates for BBH, BNS and NSBH, modeled using the SFR convolved with a time-delay distribution. For BBHs, a metallicity threshold is additionally applied. The parameters describing these models are listed in Table 1. All rates are normalized to local merger rate estimates within the 90% credible interval from GW observations, following Ref. [11].
@@ V Combining uncertainties in modeling the astrophysical GWB
To quantify the current uncertainty in the BBH background spectrum, we construct a credible band based on the ensemble of merger rate histories shown in the left panel of Fig. 11. Each merger rate history is combined with a set of hyperparameters describing the best fit models for the BBH mass and spin distributions. Specifically, we adopt the PLP model for the primary mass, a power law model for the secondary mass, and the Default spin model. The hyperparameter posterior samples are obtained via hierarchical Bayesian inference as detailed in Ref. [11]. For each sample, we compute the corresponding BBH background spectrum by evaluating Eq. 2. This involves summing contributions from a large number of BBH mergers, with masses, spins and redshifts drawn according to the population models described by the set of hyperparameters. This approach captures the combined uncertainty arising from both the merger rate evolution and the population properties of BBH systems.
@@
For BNS and NSBH systems, we estimate the uncertainty in the background in a similar manner. We combine the rates, which are presented in the middle and right panel of Fig. 11 with hyperparameters describing the mass distributions. Since these distributions remain largely unconstrained for these classes of CBCs, we consider the same simple models used for the fiducial case, with some more flexible boundaries. For BNS systems, we assume a uniform mass distribution between $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ for both components, where $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ are independently drawn from uniform distributions over the intervals $1.0,1.3]\,M_{\odot}$ and $2.0,2.5]\,M_{\odot}$, respectively. For NSBH systems, the neutron star mass distribution is taken to be the same as in the BNS case, while the black hole mass follows a log-uniform mass distribution with lower and upper bounds drawn uniformly from $3,5]\,M_{\odot}$ and $10,100]\,M_{\odot}$, respectively. Given that spin contributions are expected to have minimal impact relative to the existing uncertainties, we assume both BNS and NSBH systems to be non-spinning.
@@
Figure 12: Upper left: BBH background spectra derived from a wide ensemble of redshift evolution models, combined with uncertainties in the mass and spin distributions, based on the PLP mass model and the Default spin model. Black lines indicate the 90% credible band, providing a robust estimate of the stochastic background due to BBH mergers. Upper Right: BNS background spectra, incorporating merger rate uncertainties (as shown in Fig. 11), and a uniform mass distribution with slightly varying bounds. Black lines indicate the 90% credible band, the fiducial BNS model is shown in gray. Lower Left: NSBH background spectra, combining merger rate uncertainties with slight variations of the fiducial NSBH mass distribution. In gray we indicate the fiducial NSBH background, black lines show the 90% credible bounds. Lower Right: 90% credible bands for BBH, BNS and NSBH backgrounds. Also shown are the PI sensitivity curve for the search with data up to O3, and the target PI sensitivity curve anticipated for O5 at the $2\sigma$ level.
@@
The resulting spectra are presented in Fig. 12, which illustrates both the individual realizations and the 90% credible band for the three classes of CBCs, alongside the fiducial models introduced earlier. In the bottom right panel of Fig. 12, we show again the 90% credible bands of all CBC backgrounds, together with the $2\sigma$ power-law integrated (PI) curves [111] from the stochastic cross-correlation search following the third observing run [18], as well as the projected sensitivity for one year of coincident observation at O5 sensitivity [112, 113] with upgraded Advanced LIGO detectors (A+) [114] and Virgo detector (V+) [115]. The PI curve indicates the sensitivity to a power-law signal integrated over the entire frequency band at a given confidence level (here, $2\sigma$). If a signal lies above the PI curve, it would be detectable with at least $2\sigma$ confidence. We note that while a detection of the combined background with O4 sensitivity is unlikely, we start probing the interesting regime, if the detectors reach the sensitivity projected for O5. A detection or non-detection of a stochastic background carries important implications for population models. In the case of a non-detection, certain combinations of parameters governing the merger rate and source properties could be excluded. Nevertheless, because the background depends on numerous parameters and various modeling assumptions, isolating individual constraints remains difficult.
@@
We compare our BBH background forecast, including uncertainties, with recent results in the literature. Compared to Refs. [18, 11], we find a significantly broader uncertainty band, particularly at low frequencies. While the uncertainty in the mass distribution is incorporated in a similar way, the broader band is due to the inclusion of uncertainties in the merger rate history. In contrast, the cited works consider only the uncertainty on the local merger rate, assuming a fixed redshift evolution. Additionally, a minor difference in the spectral shape is introduced by our use of a recent waveform model. In comparison with Ref. [48], our uncertainty range is similar, even slightly narrower. This difference stems from the treatment of merger rate uncertainties: Ref. [48] varies the low-redshift power law index ($\gamma$ in Eq. 3) while keeping the high-redshift rate evolution fixed. For the background from BNS and NSBH systems, our uncertainty estimates closely match those in Ref. [11]. This agreement is expected, as the dominant source of uncertainty in these cases is the local merger rate, and we adopt similar assumptions regarding the mass distribution.
@@ VI Conclusion
We have presented a comprehensive study of the AGWB produced by CBCs, focusing on contributions from BBHs, BNSs and NSBH systems. Starting from fiducial models informed by current observational constraints, we systematically investigated how uncertainties in their mass and spin distributions, as well as different modeling choices for their merger rate histories, affect the resulting energy density spectrum. By combining the various uncertainties, we derived credible bands for the GWB from these sources, which span approximately an order of magnitude in fractional energy density. We found that while local rate uncertainties dominate the uncertainty in the GWB spectrum for BNS and NSBH systems, the merger rate history plays a more significant role for BBHs. In particular, assumptions about the time delay distribution and metallicity dependence introduce substantial variation in the predicted background.
Since the completion of this work, the GWTC-4 catalog [62] and updated population inferences [63] have been released. These updates primarily refine the underlying BBH mass model and tighten the local merger rate estimates. Their impact on the uncertainty in the predicted energy density spectra is expected to be minor, we anticipate slightly narrower uncertainty bands in Fig. 12 for BBH, BNS and NSBH. However, the dominant source of uncertainty remains the merger rate evolution. We note that the forecast in Fig. 7 of Ref. [116], which incorporates the updated population, reports even larger uncertainties for the BBH background. This difference arises because their approach estimates the low-redshift merger rate evolution directly from observed BBHs rather than using a time-delay and metallicity model, resulting in broader uncertainty compared to our modeling.
Beyond serving as a forecast for potential AGWB detection in upcoming observing runs of ground-based GW detectors, our framework provides a basis for interpreting future measurements. It enables constraints on astrophysical modeling choices and helps refining our understanding of the compact binary population across cosmic time. Future extension of this work may incorporate redshift-dependent population models and explore the impact of multiple formation channels.
@@ Acknowledgements.
We would like to thank Irina Dvorkin for helpful comments. M.E. is supported by the Swiss National Science Foundation grant Sinergia 213497 and UZH Postdoc Grant, grant no. [FK-25-109]. We thank the Physik-Institut of the University of Zurich for providing computational resources. The data that support the findings of this article are openly available at [117].

### v3

@@ IV.4 Combined merger rate uncertainty
To construct the full ensemble of merger rate histories across all classes of compact binaries, we draw random samples from the parameter space defined earlier, which governs the evolution of the merger rate. The sampling ranges and underlying distributions for these parameters are summarized in Table 1. Specifically, we vary the local rate $R_{0}$, the minimum delay time $t_{\mathrm{d,min}}$, the spectral index $\nu$ of the time-delay distribution, and exclusively for BBH systems, the metallicity threshold $Z_{\mathrm{th}}$, above which BBH formation is suppressed, and the metallicity spread $\sigma_{Z}$, which characterizes the dispersion around the mean metallicity given by Sec. IV.3. We adopt a Gaussian prior for $\nu$ (with $\mu_{\nu}=-1$ and $\sigma_{\nu}=0.25$) to reflect results from observations that a spectral index $\nu=-1$ provides a good fit to many compact binary formation models [109, 110, 59]. This choice helps to avoid a bias toward high merger rates that could arise by large negative values of $\nu$, without affecting the overall uncertainty in the combined merger rate uncertainty. It simply ensures a greater representation of merger rates consistent with a time-delay distribution with spectral index close to 1-1. While BNS formation is only weakly sensitive to metallicity, the situation is less clear for NSBH systems. Nevertheless, given the already broad uncertainties in the NSBH merger rate estimates, we choose not to introduce additional uncertainty from metallicity effects. The resulting CBC merger rates are shown in Fig. 11. While the local BBH merger rate is relatively well constrained by direct observations, the rate at redshifts $z>2$ can vary by more than two orders of magnitude depending on the model assumptions. The local rates of BNS and NSBH coalescences are much less constrained.
@@
Figure 11: Ensembles of merger rates for BBH, BNS, and NSBH, modeled using the SFR convolved with a time-delay distribution. For BBHs, a metallicity threshold is additionally applied. The parameters describing these models are listed in Table 1. All rates are normalized to local merger rate estimates within the 90% credible interval from GW observations, following Ref. [11].
@@ V Combining uncertainties in modeling the astrophysical GWB
To quantify the current uncertainty in the BBH background spectrum, we construct a credible band based on the ensemble of merger rate histories shown in the left panel of Fig. 11. Each merger rate history is combined with a set of hyperparameters describing the best-fit models for the BBH mass and spin distributions. Specifically, we adopt the PLP model for the primary mass, a power-law model for the secondary mass, and the default spin model. The hyperparameter posterior samples are obtained via hierarchical Bayesian inference, as detailed in Ref. [11]. For each sample, we compute the corresponding BBH background spectrum by evaluating Eq. 2. This involves summing contributions from a large number of BBH mergers, with masses, spins, and redshifts drawn according to the population models described by the set of hyperparameters. This approach captures the combined uncertainty arising from both the merger rate evolution and the population properties of BBH systems.
@@
For BNS and NSBH systems, we estimate the uncertainty in the background in a similar manner. We combine the rates, which are presented in the middle and right panels of Fig. 11, with hyperparameters describing the mass distributions. Since these distributions remain largely unconstrained for these classes of CBCs, we consider the same simple models used for the fiducial case, with some more flexible boundaries. For BNS systems, we assume a uniform mass distribution between $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ for both components, where $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ are independently drawn from uniform distributions over the intervals $1.0,1.3]\,M_{\odot}$ and $2.0,2.5]\,M_{\odot}$, respectively. For NSBH systems, the neutron star mass distribution is taken to be the same as in the BNS case, while the black hole mass follows a log-uniform mass distribution with lower and upper bounds drawn uniformly from $3,5]\,M_{\odot}$ and $10,100]\,M_{\odot}$, respectively. Given that spin contributions are expected to have minimal impact relative to the existing uncertainties, we assume both BNS and NSBH systems to be nonspinning.
@@
Figure 12: Upper left: BBH background spectra derived from a wide ensemble of redshift evolution models, combined with uncertainties in the mass and spin distributions, based on the PLP mass model and the default spin model. Black lines indicate the 90% credible band, providing a robust estimate of the stochastic background due to BBH mergers. Upper Right: BNS background spectra, incorporating merger rate uncertainties (as shown in Fig. 11), and a uniform mass distribution with slightly varying bounds. Black lines indicate the 90% credible band, and the fiducial BNS model is shown in gray. Lower Left: NSBH background spectra, combining merger rate uncertainties with slight variations of the fiducial NSBH mass distribution. In gray we indicate the fiducial NSBH background, and black lines show the 90% credible bounds. Lower Right: 90% credible bands for BBH, BNS, and NSBH backgrounds. Also shown are the PI sensitivity curve for the search with data up to O3, and the target PI sensitivity curve anticipated for O5 at the $2\sigma$ level.
@@
The resulting spectra are presented in Fig. 12, which illustrates both the individual realizations and the 90% credible band for the three classes of CBCs, alongside the fiducial models introduced earlier. In the bottom right panel of Fig. 12, we show again the 90% credible bands of all CBC backgrounds, together with the $2\sigma$ power-law integrated (PI) curves [111] from the stochastic cross-correlation search following the third observing run [18], as well as the projected sensitivity for one year of coincident observation at O5 sensitivity [112, 113] with upgraded Advanced LIGO detectors (A+) [114] and the Virgo detector (V+) [115]. The PI curve indicates the sensitivity to a power-law signal integrated over the entire frequency band at a given confidence level (here, $2\sigma$). If a signal lies above the PI curve, it would be detectable with at least $2\sigma$ confidence. We note that while a detection of the combined background with O4 sensitivity is unlikely, we start probing the interesting regime if the detectors reach the sensitivity projected for O5. A detection or nondetection of a stochastic background carries important implications for population models. In the case of a nondetection, certain combinations of parameters governing the merger rate and source properties could be excluded. Nevertheless, because the background depends on numerous parameters and various modeling assumptions, isolating individual constraints remains difficult.
@@
We compare our BBH background forecast, including uncertainties, with recent results in the literature. Compared to Refs. [18, 11], we find a significantly broader uncertainty band, particularly at low frequencies. While the uncertainty in the mass distribution is incorporated in a similar way, the broader band is due to the inclusion of uncertainties in the merger rate history. In contrast, the cited works consider only the uncertainty on the local merger rate, assuming a fixed redshift evolution. Additionally, a minor difference in the spectral shape is introduced by our use of a recent waveform model. In comparison with Ref. [48], our uncertainty range is similar, even slightly narrower. This difference stems from the treatment of merger rate uncertainties: Ref. [48] varies the low-redshift power-law index $\gamma$ in Eq. 3] while keeping the high-redshift rate evolution fixed. For the background from BNS and NSBH systems, our uncertainty estimates closely match those in Ref. [11]. This agreement is expected, as the dominant source of uncertainty in these cases is the local merger rate, and we adopt similar assumptions regarding the mass distribution.
@@ VI Conclusion
We have presented a comprehensive study of the AGWB produced by CBCs, focusing on contributions from BBH, BNS, and NSBH systems. Starting from fiducial models informed by current observational constraints, we systematically investigated how uncertainties in their mass and spin distributions, as well as different modeling choices for their merger rate histories, affect the resulting energy density spectrum. By combining the various uncertainties, we derived credible bands for the GWB from these sources, which span approximately an order of magnitude in fractional energy density. We found that while local rate uncertainties dominate the uncertainty in the GWB spectrum for BNS and NSBH systems, the merger rate history plays a more significant role for BBHs. In particular, assumptions about the time-delay distribution and metallicity dependence introduce substantial variation in the predicted background.
Since the completion of this work, the GWTC-4 catalog [62] and updated population inferences [63] have been released. These updates primarily refine the underlying BBH mass model and tighten the local merger rate estimates. Their impact on the uncertainty in the predicted energy density spectra is expected to be minor—we anticipate slightly narrower uncertainty bands in Fig. 12 for BBH, BNS, and NSBH systems. However, the dominant source of uncertainty remains the merger rate evolution. We note that the forecast in Fig. 7 of Ref. [116], which incorporates the updated population, reports even larger uncertainties for the BBH background. This difference arises because their approach estimates the low-redshift merger rate evolution directly from observed BBHs rather than using a time-delay and metallicity model, resulting in broader uncertainty compared to our modeling.
Beyond serving as a forecast for potential AGWB detection in upcoming observing runs of ground-based GW detectors, our framework provides a basis for interpreting future measurements. It enables constraints on astrophysical modeling choices and helps in refining our understanding of the compact binary population across cosmic time. Future extension of this work may incorporate redshift-dependent population models and explore the impact of multiple formation channels.
@@ Acknowledgements.
We would like to thank Irina Dvorkin for helpful comments. M.E. is supported by the Swiss National Science Foundation Grant Sinergia 213497 and UZH Postdoc Grant, No. [FK-25-109]. We thank the Physik-Institut of the University of Zurich for providing computational resources. The data that support the findings of this article are openly available at [117].